OpenStudy (anonymous):

I WILL MEDAL AND FAN a+ 2b = 10 2a + b =6 Solve this system of equations for a and b. A) a = 102/5 , b = 13 B) a = 14/3 , b = 8/3 C) a = 2/3 , b = 14/3 D) a = - 2/5 , b = 34/5

rishavraj (rishavraj):

how u wanna solve : by elimination or by substitution ??

OpenStudy (anonymous):

Elimination

rishavraj (rishavraj):

multiply 2nd equation by -2 and add it with 1st equation ...

OpenStudy (anonymous):

I really do apologize but this math subject kinda confuses me? ):

rishavraj (rishavraj):

like it would be : a + 2b = 10 and -4a - 2b = -12 now add both equations... b gets eliminated

OpenStudy (anonymous):

alright!

rishavraj (rishavraj):

:)))))

OpenStudy (anonymous):

"A" must be eliminated too?

OpenStudy (anonymous):

Because i tried solving it and i dont think A was right?

rishavraj (rishavraj):

wht do u mean???

OpenStudy (anonymous):

A is wrong right?

rishavraj (rishavraj):

yup.....lol stop guessing if u are... :P after b gets eliminated....... -3a = -2 so 3a = 2 a = 2/3 now get b

OpenStudy (anonymous):

thanks i got it now though!

rishavraj (rishavraj):

lol all bcoz a = 2/3 is only in one option